OBSERVATION AND ANALYSIS OF BOUND → FREE TRANSITION OF ScI_{3} PHOTOEXCITED IN THE ULTRAVIOLET

\underline{LOANN D. POMMAREL}, J. DARBY HEWITT, THOMAS C. GALVIN, and J. GARY EDEN, Laboratory for Optical Physics and Enginnering, Department of Electrical and Computer Engineering, University of Illinois at Urbana-Champaign, 607 East Healey Street, Champaign, IL 61820.

Bound to free transitions of ScI were observed for the first time after photoexcitation of ScI_{3} in the ultraviolet by 193 nm radiation from an ArF excimer laser. Broad fluorescence features typical of molecular transitions between different vibrational energy levels of Sc*+I have been measured in the near-infrared between 860 and 940 nm. Many narrow spectral lines have also been observed, but were attributed to fluorescence of atomic Scandium or Iodine. Experimental and simulation results will be compared and presented.